OpenStudy (anonymous):

If 74.0 sodium is mixed with 27.0 nitrogen gas, what mass of sodium nitride forms?

OpenStudy (anonymous):

6Na + N2 = 2Na3N Na will be the limiting factor. M.W of NA = 23 74/23 = 3.22 moles 3.218 /6 = .536 moles M.W of N = 14 27/14 = 1.93 moles /2 = .964 moles 3.22 moles Na * (2 moles Na3N) / 6 Moles of NA = 1.07 moles of Na3N formed. 1.07 moles Na3N * 83 (M.W of Na3N) = 88.71 grams of Na3N formed. WOW that took forever :(

OpenStudy (anonymous):

Thank You some much :D

OpenStudy (anonymous):

If the reaction in has a percent yield of 69.0 %, how much sodium nitride is actually produced?